# Using MultiPrecisonArrays.jl: Iterative Refinement in Julia

### v7

@@ Abstract
This document is for  v0.1.3  of the package. We used both the AppleAccelerate [10] framework and openBLAS for the examples and the timings may be different from the older versions. All the computations in this paper were done on an Apple Mac Mini with an M2 Pro processor and 8 performance cores. The examples were done with several different versions of MultiPrecisionArrays.jl , Julia, and the OS. Hence the examples are indicative of the performance of IR, but the reader’s results may differ in timings and iteration counts. The qualitative results are the same.
@@
Interative Refinement, Mixed-Precision Arithmetic, Interprecision Transfers, Julia
@@ 1 Introduction
IR is a perfect example of a storage/time tradeoff. To solve a linear system ${\bf A}{\bf x}={\bf b}$ in $R^{N}$ with IR, one incurs the storage penalty of making a low precision copy of ${\bf A}$ and reaps the benefit of only having to factor the low precision copy.
In most of this paper we consider IR using two precisions. We use the naming convention of [1, 3, 4, 6] for these precisions and the others we will consider later. We start with the case where we store the data ${\bf A}$ and ${\bf b}$ in the higher or working precision and compute the factorization in the lower or factorization preicsion. Following standard Julia type notation, we will let TW and TF be the working and factorization precision types. So, for example
@@
In a typical use case, TW will be double and TF will be single. We will make precision and inter precision transfers explicit in our algorithmic descriptions.
@@
One must be clear on the meanings of ‘‘factor in low precision’’ and ${\bf d}={\bf U}^{-1}{\bf L}^{-1}{\bf r}$ to implement the algorithm. As we indicated above, the only way to factor ${\bf A}$ in low precision is to make a copy and factor that copy. We must introduce some notation for that. We let ${\cal F}_{p}$ be the set of floating point numbers in precision $p$, $u_{p}$ the unit roundoff in that precision, and $fl_{p}$ the rounding operator. Similarly we let ${\cal F}^{N}_{p}$, ${\cal F}^{N\times N}_{p}$ denote the vectors and matrices in precision $p$. We let $I_{p}^{q}$ denote the copying operator from precision $p$ to precision $q$. When we are not explicitly specifying the precisions, we will use $W$ and $F$ as sub and superscripts for the working and factorization precisions. When we are discussing specific use cases out sub and superscripts will the $d$, $s$, and $h$ for double, single, and half precision.
So, factoring a high-precision matrix ${\bf A}\in{\cal F}^{N\times N}_{H}$ in low precision $L$ means copy ${\bf A}$ into low precision and obtain
@@
and then factor ${\bf A}_{L}={\bf L}{\bf U}$.
The current version of MultiPrecisionArrays.jl  v0.1.3  requires that the type of ${\bf A}$ be
@@ 2 Integral Equations Example
The submodule MultiPrecisionArrays.Examples has an example which we will use repeatedly. The function Gmat(N) returns the $N$ point trapezoid rule discretization of the Greens operator for $-d^{2}/dx^{2}$ on $[0,1]$
[…]
@@
In the examples we will use Gmat to build a matrix ${\bf A}=I-\alpha{\bf G}$. In the examples we will use $\alpha=1.0$, a very well conditioned case, and $\alpha=800.0$. This latter case is very near singularity.
2.1 Classic Example: Double-Single Precision
While MultiPrecisionArrays.jl was designed for research, it is useful in applications in the classic case where the working precision is double and the factorization precision is single. This case avoids the (very interesting) problems with half precision.
@@ 2.2 Running MultiprecisionArrays: I
The most simple way to use this package is to combine the constrution of the MPArray with the factorization of the low precision copy of ${\bf A}$. One does this with the mplu command.
@@
It is no surprise that the factorization in single precision took roughly half as long as the one in double. In the double-single precision case, iterative refinement is a great expample of a time/storage tradeoff. You have to store a low precision copy of , so the storage burden increases by 50% and the factoriztion time is cut in half. The advantages of IR increase as the dimension increases. IR is less impressive for smaller problems and can even be slower
@@
[…]
@@
2.3 Harvesting Iteration Statistics: Part I
You can get some iteration statistics by using the reporting keyword argument to the solvers. The easiest way to do this is with the backslash command. When you use this option you get a data structure with the solution and the residual history.
@@ 2.3 Harvesting Iteration Statistics: Part I
9.99878e-01
1.21892e-04
5.25805e-11
2.56462e-14
1.33227e-15
@@
5-element Vector{Float64}:
9.99878e-01
6.17721e-07
3.84581e-13
7.99361e-15
8.88178e-16
So the second iteration is much better, but the iteration terminated after four iterations in both cases.
@@
9-element Vector{Float64}:
9.99878e-01
4.58739e-03
1.86362e-05
7.36240e-08
2.89855e-10
1.14420e-12
3.44169e-14
3.10862e-15
6.66134e-16
@@ 2.4 Options and data structures for mplu
The function mplu has four keyword arguments. The easy one to understand is TF which is the precision of the factoriztion. Julia has support for single (Float32) and half (Float16) precisions. If you set TF=Float16 then low precision will be half. Don’t do that unless you know what you’re doing. Using half precision is a fast way to get incorrect results. Look at § 3 for a bit more bad news.
TR is the residual precision. The MPArray structure preallocates storage for the residual and a local copy of the solution in precision TR. For most applications (but see § 6) TR = TW.
@@
the LU factors are in low precision and the residual $r$ is in high precision. If you let Julia and LAPACK figure out what to do, then the solves will be done in high precision and the entries in the LU factors will be comverted to high precision with each binary operation. The output $d$ will be in high precision. This is called interprecision transfer on-the-fly and onthefly = true will tell the solvers to do it that way. You have $N^{2}$ interprecsion transfers with each solve and, as we will see, that can have a non-trivial cost.
@@
When you set onthefly = false the solve converts $r$ to low precision, does the solve entirely in low precision, and then promotes $d$ to high precision. This is in-place interprecision transfer. You need to be careful to avoid overflow and, more importantly, underflow when you do that and we scale $r$ to be a unit vector before conversion to low precisiion and reverse the scaling when we promote $d$. We take care of this for you.
@@
mplu calls the constructor for the multiprecision array and then factors the low precision matrix. In some cases, such as nonlinear solvers, you will want to separate the constructor and the factorization. I will provide code for this in a future release. I do not export the constructor and if you mess with it, remember that mplu! overwrites the low precision copy of A with the factors. The factoriztion object is different from the mulitprecision array, even though they share storage. Be careful with this.
@@ 3 Half Precision
At this point Julia has no native support for BFloat16. Progress is being made and we expect to suppport BFloat16 in the future.
Doing the factorization in half precision (Float16 or BFloat16) will not speed up the solver, in fact it will make the solver slower. The reason for this is that LAPACK and the BLAS do not (YET [5]) support half precision, so all the clever stuff in there is missing. We provide a half precision LU factorization for Float16 /src/Factorizations/hlu!.jl that is better than nothing. It’s a hack of Julia’s generic_lu! with threading and a couple compiler directives. Even so, it’s 2 -- 5 times slower than a double precision LU. Half precision support is coming [5] and Julia and Apple support Float16 in hardware. Apple also has hardware support for BFloat16. However, for now, at least for desktop computing, half precision is for research in iterative refinement, not applications.
@@
N Double Single Half Ratio 1024 3.96e-03 2.80e-03 2.48e-02 6.26e+00 2048 2.23e-02 1.43e-02 6.91e-02 3.09e+00 4096 1.54e-01 8.51e-02 3.08e-01 1.99e+00 8192 1.14e+00 6.02e-01 4.40e+00 3.87e+00
@@ 5.1 Terminating the while loop
(3) $\|{\bf r}\|<C_{e}u_{r}(\|{\bf b}\|+\|{\bf A}\|\|{\bf x}\|).$
The prefactors $C_{r}$ and $C_{e}$ are algorithmic paramters. In the current version of MultiPrecisionArrays.jl the are not adjustable by the user. The current values, which I may change, are
$C_{e}=1.0\mbox{ and }C_{r}=20.0$
The value of $C_{e}$ is consistent with the recommendations in [8].
@@
Termination on small residual is the default because computing $\|{\bf A}\|$ is $N^{2}$ work is more expensive that a few IR iterations. I am using $\|{\bf A}\|_{1}$ because that takes less time (data locality) than $\|{\bf A}\|_{\infty}$ but it is still a pain. The reason $C_{r}>C_{e}$ is that putting $\|A\|$ in the termination criterion is very useful and making $C_{r}$ large helps compensate for that.
The problem with these criteria is that IR can stagnate, especially for ill-conditioned problems, before the termination criterion is attained. We detect stagnation by looking for a unacceptable decrease (or increase) in the residual norm. So we will also terminate the iteration if

### v8

@@ Abstract
This document is for  v0.1.4  of the package. We used both the AppleAccelerate [10] framework and openBLAS for the examples and the timings may be different from the older versions. All the computations in this paper were done on an Apple Mac Mini with an M2 Pro processor and 8 performance cores. The examples were done with several different versions of MultiPrecisionArrays.jl , Julia, and the OS. Hence the examples are indicative of the performance of IR, but the reader’s results may differ in timings and iteration counts. The qualitative results are the same.
@@
Iterative Refinement, Mixed-Precision Arithmetic, Interprecision Transfers, Julia
@@ 1 Introduction
We focus on two views of IR. One is as a perfect example of a storage/time tradeoff. To solve a linear system ${\bf A}{\bf x}={\bf b}$ in $R^{N}$ with IR, one incurs the storage penalty of making a low precision copy of $\bf A$} and reaps the benefit of only having to factor the low precision copy.
The other viewpoint is as a way to address very ill-conditioned problems, which was the purpose of the original paper on the topic [21]. Here one stores and factors $\bf A$} in one precision but evaluates the residual in a higher precision.
In most of this paper we consider IR using two precisions. We use the naming convention of [1, 3, 4, 6] for these precisions and the others we will consider later. We start with the case where we store the data $\bf A$} and $\bf b$} in the higher or working precision and compute the factorization in the lower or factorization precision. Following standard Julia type notation, we will let $TW$ and $TF$ be the working and factorization precision types. So, for example
@@
In our first case, for example, $TW$ will be double and $TF$ will be single. In the second case, $TW$ and $TF$ will both be single precision and we will compute the residual in precision $TR$, which will be double. We will make precision and inter precision transfers explicit in our algorithmic descriptions.
@@
One must be clear on the meanings of ‘‘factor in precision $TF$’’ and ${\bf d}={\bf U}^{-1}{\bf L}^{-1}{\bf r}$ to implement the algorithm. As we indicated above, if $TW$ is a higher precision than $TF$, the only way to factor $\bf A$} in low precision is to make a copy and factor that copy. We must introduce some notation for that. We let ${\cal F}_{p}$ be the set of floating point numbers in precision $p$, $u_{p}$ the unit roundoff in that precision, and $fl_{p}$ the rounding operator. Similarly we let ${\cal F}^{N}_{p}$, ${\cal F}^{N\times N}_{p}$ denote the vectors and matrices in precision $p$. We let $I_{p}^{q}$ denote the copying operator from precision $p$ to precision $q$. When we are not explicitly specifying the precisions, we will use $W$ and $F$ as sub and superscripts for the working and factorization precisions. When we are discussing specific use cases out sub and superscripts will the $d$, $s$, and $h$ for double, single, and half precision.
So, factoring a high-precision matrix ${\bf A}\in{\cal F}^{N\times N}_{H}$ in low precision $TF$ means copy $\bf A$} into low precision and obtain
@@
and then factor ${\bf A}_{F}={\bf L}{\bf U}$.
The current version of MultiPrecisionArrays.jl  v0.1.4  requires that the type of $\bf A$} be
@@ 2 Integral Equations Example
The submodule MultiPrecisionArrays.Examples has an example which we will use repeatedly. The function Gmat(N) returns the $N$ point trapezoid rule discretization of the Greens operator for $d^{2}/dx^{2}$ on $[0,1]$
$Gu(x)=\int_{0}^{1}g(x,y)u(y)\,dy$
@@
In the examples we will use Gmat to build a matrix ${\bf A}=I-\alpha{\bf G}$. In the examples we will use $\alpha=1.0$, a very well conditioned case, $\alpha=800.0$, a somewhat ill-conditioned case, and $\alpha=799.0$, a very ill-conditioned case. The latter case is very near singularity.
2.1 Two Classic Examples
While MultiPrecisionArrays.jl was designed for research, it is useful in applications. We present two examples. The first is one where $TW=TR$ is double and $TF$ is single. This example illustrates the essential ideas in IR and shows how to use MultiPrecisionArrays.jl in the most direct way. We follow this example with an illustration on how one harvests iteration statistics.
The second example is more subtle. Here $TR$ is double, $TF=TW$ is single, and the matrix is highly ill-conditioned. Here the termination criteria is different and we look at more iteration statistics.
2.1.1 Double-Single Precision
In this case the working precision is double and the factorization precision is single. This case avoids the (very interesting) problems with half precision.
@@ 2.2 Running MultiprecisionArrays: I
The most simple way to use this package is to combine the construction of the MPArray with the factorization of the low precision copy of $\bf A$}. One does this with the mplu command.
@@
It is no surprise that the factorization in single precision took roughly half as long as the one in double. In the double-single precision case, iterative refinement is a great example of a time/storage tradeoff. You have to store a low precision copy of , so the storage burden increases by 50% and the factorization time is cut in half. The advantages of IR increase as the dimension increases. IR is less impressive for smaller problems and can even be slower
@@
Now for the solves. Both lu and mplu produce a Julia factorization object and $\backslash$ works with both. You have to be a bit careful because MPA and A share storage. So I will use lu instead of lu! when factoring $\bf A$}.
@@
2.2.1 Harvesting Iteration Statistics: Part I
You can get some iteration statistics by using the reporting keyword argument to the solvers. The easiest way to do this is with the backslash command. When you use this option you get a data structure with the solution, the residual norm history, and the history of the correction norms.
@@ 2.2.1 Harvesting Iteration Statistics: Part I
6-element Vector{Float64}:
1.00000e+00
1.64859e-05
1.02649e-10
6.35048e-14
4.44089e-16
2.22045e-16
julia> # The history of the correction norms is shorter.
julia> mpout.dhist
@@
1.00002e+00
1.62124e-05
1.02649e-10
6.35100e-14
4.45650e-16
One computes a correction after all resduals but the last, when the iteration terminates on small resdiual norm.
@@
6-element Vector{Float64}:
1.00000e+00
3.10192e-06
1.04403e-11
6.13953e-14
4.44089e-16
2.22045e-16
There is very little difference in this case.
@@
1.00000e+00
4.59118e-03
1.86491e-05
7.36737e-08
2.90049e-10
1.15497e-12
9.61453e-14
1.33227e-15
2.22045e-16
In § 2.2.2 we consider a much more ill-conditioned example.
2.2.2 Evaluating the residual in extended precision
IR was invented for this case and the idea comes from [21]. I am using the notation from [6, 1]. The idea is to evaluate the residual in a precision TR higher than the working precision. One does this by promoting the solution to $TR$ and then using the fact that if $TR$ is a higher precision than $TW$ and $\circ$ is any of $+,\times,-,\div$, $a\in{\cal F}_{R}$ and $b\in{\cal F}_{W}$ that
(1) $fl_{R}(a\circ b)=fl_{R}(a\circ I_{W}^{R}(b)).$
This implies that
${\bf r}={\bf b}-{\bf A}(I_{W}^{R}{\bf x})=(I_{W}^{R}{\bf b})-(I_{W}^{R}{\bf A})(I_{W}^{R}{\bf x}).$
After the computation of $\bf r$} in precision TR one must downcast it to TW before doing the solve for the correction. You must do interprecision transfers on the fly in this case. Since the convergence is based on the norm of the residual in precision TR, you are really solving a promoted problem [15]
$(I_{W}^{R}A)x=I_{W}^{R}b$
and, by driving the residual to a small value can mitigate ill-conditioning to some degree. MultiPrecisionArrays.jl allows you to do this with the multiprecision factorization you get from mplu. The Krylov-IR solvers can also converge even if the factorization is not accurate enough to get convergence of IR (see § 2.3).
The classic example is to let TW and TF be single precision and TR be double. The storage penalty is that you must store two copies of $A$, one for the residual computation and the other for the factorization. Here is an example with a badly conditioned matrix. You must tell mplu to factor in the working precision and then use the kwargs in the solver to set TR.
The example in this section is highly ill-conditioned. We base the example on the integral equation
(2) $u-799.0\ Gu=f(x)\equiv 1-\frac{799x(1-x)}{2}.$
The solution is $u\equiv 1$.
julia> using MultiPrecisionArrays
julia> using MultiPrecisionArrays.Examples
julia> N=4096; alpha=799.0; AD=I - alpha*Gmat(N);
# conditioning is bad
julia> cond(AD,Inf)
2.35899e+05
# Set up the single precision computation
# and use the right side from the integral equation
julia> h=1.0/(N-1); x=collect(0:h:1); bd = 1.0 .- alpha*x.*(1.0 .- x)*.5;
# Solving in double gives the accuracy you'd expect
[…]
julia> norm(u .- 1.0)
3.16529e-10
# Now move the problem to single precision
julia> A = Float32.(AD); xe=ones(Float32,N); b=Float32.(bd)
julia> # You can see the ill-conditioning
[…]
1.37073e-02
julia> # The solution of the promoted problem is better.
[…]
1.44238e-04
julia> # Make sure TF is what it needs to be for this example
julia> # Set TR in the call to mplu.
julia> AF = mplu(A; TF=Float32, TR=Float64);
julia> # Use the multiprecision array to solve the problem.
julia> mrout = \(AF, b; reporting=true);
julia> # look at the residual history
julia> mrout.rhist
6-element Vector{Float64}:
9.88750e+01
6.38567e-03
1.24560e-06
9.55874e-08
8.49478e-08
8.49478e-08
julia> # Compare the solution to the solution of the promoted problem
julia> norm(mrout.sol - xp,Inf)
5.95629e-08
julia> # That's consistent with theory.
juila> # So the solution is ok?
julia> norm(mrout.sol - xe, Inf)
1.44243e-04
julia> # Yes.
AF = mplu(A; TF=Float32, TR=Float64);
with
AF = mplu(A; TR=Float64);
The iteration fails to converge because $AF$ is not accurate enough.
julia> AF = mplu(A; TR=Float64);
julia> mout16=\(AF, b; reporting=true);
julia> mout16.rhist
4-element Vector{Float64}:
9.88750e+01
3.92451e+00
3.34292e-01
2.02204e-01
2.3 IR-Krylov
MultiPrecisionArrays.jl supports IR-GMRES and IR-BiCGSTAB for TR > TW. You get this to work just like in mplu by using the keyword argument TR. We will continue with the example in this section and do that. For this example the default basis size of $10$ is not enough, so we use 20.
julia> GF = mpglu(A; TR=Float64, basissize=20);
julia> moutG = \(GF, b; reporting=true);
julia> moutG.rhist
8-element Vector{Float64}:
9.88750e+01
7.59075e-03
1.48842e-05
2.17281e-07
8.60429e-08
7.45077e-08
7.91866e-08
7.53089e-08
julia> moutG.dhist
7-element Vector{Float32}:
1.03306e+00
3.23734e-02
2.83073e-04
8.92629e-06
1.55432e-07
6.05685e-08
5.96633e-08
# We got the solution to the promoted problem...
[…]
julia> norm(xp-moutG.sol, Inf)
5.95825e-08
# IR-BiCGSTAB would take fewer iterations than IR-GMRES had we used
# the default basis size because there's no storage
julia> BF = mpblu(A; TR=Float64);
julia> moutB = \(BF, b; reporting=true);
julia> moutB.rhist
5-element Vector{Float64}:
9.88750e+01
1.86437e-07
7.53089e-08
7.53089e-08
7.53089e-08
julia> moutB.dhist
4-element Vector{Float32}:
1.00227e+00
2.05457e-06
5.95821e-08
5.95821e-08
julia> norm(xp - moutB.sol, Inf)
5.95825e-08
@@ 2.4 Options and data structures for mplu
The function mplu has four keyword arguments. The easy one to understand is TF which is the precision of the factorization. Julia has support for single (Float32) and half (Float16) precisions. If you set TF=Float16 then low precision will be half. Don’t do that unless you know what you’re doing. Using half precision is a fast way to get incorrect results. Look at § 3 for a bit more bad news.
TR is the residual precision. The MPArray structure preallocates storage for the residual and a local copy of the solution in precision TR. For most applications (but see § 2.2.2) TR = TW.
@@
the LU factors are in low precision and the residual $r$ is in high precision. If you let Julia and LAPACK figure out what to do, then the solves will be done in high precision and the entries in the LU factors will be converted to high precision with each binary operation. The output $d$ will be in high precision. This is called interprecision transfer on-the-fly and onthefly = true will tell the solvers to do it that way. You have $N^{2}$ interprecision transfers with each solve and, as we will see, that can have a non-trivial cost.
@@
When you set onthefly = false the solve converts $r$ to low precision, does the solve entirely in low precision, and then promotes $d$ to high precision. This is in-place interprecision transfer. You need to be careful to avoid overflow and, more importantly, underflow when you do that and we scale $r$ to be a unit vector before conversion to low precision and reverse the scaling when we promote $d$. We take care of this for you.
@@
mplu calls the constructor for the multiprecision array and then factors the low precision matrix. In some cases, such as nonlinear solvers, you will want to separate the constructor and the factorization. I will provide code for this in a future release. I do not export the constructor and if you mess with it, remember that mplu! overwrites the low precision copy of A with the factors. The factorization object is different from the multiprecision array, even though they share storage. Be careful with this.
@@ 3 Half Precision
At this point Julia has no native support for BFloat16. Progress is being made and we expect to support BFloat16 in the future.
Doing the factorization in half precision (Float16 or BFloat16) will not speed up the solver, in fact it will make the solver slower. The reason for this is that LAPACK and the BLAS do not (YET [5]) support half precision, so all the clever stuff in there is missing. We provide a half precision LU factorization for Float16 /src/Factorizations/hlu!.jl that is better than nothing. It’s a hack of Julia’s generic_lu! with threading and a couple compiler directives. Even so, it’s 3 -- 7 times slower than a double precision LU. Half precision support is coming [5] and Julia and Apple support Float16 in hardware. Apple also has hardware support for BFloat16. However, for now, at least for desktop computing, half precision is for research in iterative refinement, not applications.
@@
N Double Single Half Ratio 512 1.37e-03 6.32e-04 3.84e-03 2.82e+00 1024 4.97e-03 2.82e-03 3.51e-02 7.07e+00 2048 2.67e-02 1.49e-02 1.91e-01 7.15e+00 4096 1.64e-01 8.32e-02 7.58e-01 4.63e+00 8192 1.25e+00 7.18e-01 5.67e+00 4.53e+00
@@ 5.1 Terminating the while loop
There are many parameters in the termination criteria and the defaults in MultiPrecisionArrays.jl are reasonable for most applications. These defaults differ slightly from the recommendations in [8] and we explain that later in this section.
@@
(5) $\|{\bf r}\|<C_{e}u_{w}(\|{\bf b}\|+\|{\bf A}\|\|{\bf x}\|).$
You make the choice between (4) and (5) when you compute the multiprecision factorization.
The prefactors $C_{r}$ and $C_{e}$ are algorithmic parameters $C_{r}=1.0$ is the default in MultiPrecisionArrays.jl . The recommendation in [8] is that $C_{e}=1.0$, which is the default in MultiPrecisionArrays.jl .
If TR is wider than TW, then I assume that you are solving a highly ill-conditioned problem and terminate when the norms of the corrections stagnate
(6) $\|{\bf d}_{new}\|\geq R_{max}\|{\bf d}_{old}\|.$
This is a bit different from the approach in [6] and I may change to that way in the future. Using (6) is simpler, but might take one more iteration to drive the error to $u_{w}$, as the theory predicts if IR converges.
5.1.1 Residual or Backward Error
Using the normwise backward error is more costly because the computation of $\|{\bf A}\|$ is $N^{2}$ work and is more expensive that a few IR iterations.
In MultiPrecisionArrays.jl we do this when we compute $\|{\bf A}\|$ during the multiprecision factorization. After we copy $\bf A$} to the factorization precision, we can take the norm of the low-precision copy before we factor it. Specifically, we compute $\|{\bf A}\|_{1}$ in the factorization precision if TF = Float32 or TF = Float64. If TF is half precision, which is still slow in our desktop environment, we compute $\|{\bf A}\|$ in the working precision TW. We control this via the residterm kwarg for the factorizations.
@@ 5.1.1 Residual or Backward Error
5.1.2 Protection Against Stagnation
The problem with either of (4) or (5) is that IR can stagnate, especially for ill-conditioned problems, before the termination criterion is attained. We detect stagnation by looking for a unacceptable decrease (or increase) in the residual norm. So we will terminate the iteration if